# Propagation speeds of relativistic conformal fluids from a generalized relaxation time approximation

### v1

@@ IV.2 Second order
The constraint eq. (22) yields
${d^{\prime}}^{\mu}=\frac{\left(1\right)}{5t^{3}}D_{3}\sigma^{\mu\rho}_{,\rho}$ (82)
with
We observe that $D_{3}$ vanishes when the Anderson Witting collision term is chosen.
This concludes the construction of the second order Chapman-Enskog solution. We now must use $\chi_{1}$ and $\chi_{2}$ as templates whereby to identify which functions of momentum to include into a general parameterization of $\chi$.
V Hydrodynamics from the moments approach
Assuming a parameterization of the form eq. (2) begs the question of which functions $X^{\alpha}$ ought to be included. In search of guidance, we look at the functions of momentum which actually show up in the second order Chapman-Enskog solution. From eqs. (67) and (81) we see that the second order $\chi$ may be regarded as a linear combination of four (tensor) functions
$X^{1}$ $=$ $\frac{1}{2t^{2}F}\Lambda^{\mu\rho}_{\lambda\tau}p^{\lambda}p^{\tau}$
where
$\gamma=\frac{\Theta^{3}_{6}}{\Theta^{2}_{5}}$ (85)
$\Lambda^{\mu\nu\rho}_{\lambda\sigma\tau}$ is the projection over transverse, totally symmetric and traceless third order tensors. Thus the $X^{\alpha}$ tensors are totally symmetric, traceless and transverse with respect to the Landau-Lifshitz velocity, and moreover they obey the orthogonality condition
We write the Grad 1pdf as
$f=e^{\sum_{\alpha=0}^{4}C_{\alpha}X^{\alpha}}$ (87)
Where $X^{0}=p^{\mu}$ and $C_{0}=\beta_{\mu}$, and the $C_{\alpha}$ with $\alpha=1-4$ are totally symmetric, traceless and transverse tensors. The equations of motion are the moment equations
$\int\;Dp\;X^{\alpha}\left[p^{\mu}f_{,\mu}-f_{0}I_{coll}\left[\chi\right]\right]=0$ (88)
where $\chi=\sum_{\alpha=1}^{4}C_{\alpha}X^{\alpha}$, which is consistent with the constraint eq. (22), and $f_{0}=e^{\beta_{\mu}p^{\mu}}$. The equation corresponding to $\alpha=0$ is just energy-momentum conservation.
VI Propagation speeds
We are now ready to derive the propagation speeds $v$ for linearized fluctuations around an equilibrium solution $\beta_{\mu}=\beta_{0\mu}=$ constant, $C^{\alpha}=0$, $\alpha=1-4$. To obtain them, we shall use that the propagation speeds $v$ may be derived from the dispersion relations in the free streaming limit, which take the form $\omega=vk$, as we shall show presently.
Observe that taking the free streaming limit does not make the choice of the function $F$ in the collision term irrelevant, because the tensors $X^{\alpha}$ which make up $\chi$ depend on it, see eqs. (84). That is, we have written a parameterized 1-pdf general enough to include a second order Chapman-Enskog solution of the collisional theory, and now we take the free streaming limit to investigate the propagation speeds for fronts. We shall find that the propagation speeds depend on the choice of $F$, and for a wide range of choices, are maximized by the Anderson-Witting ansatz $F[x]=x$.
VI.1 Propagation speeds and free streaming limit
The last term comes from the collision integral.
Let us seek a solution representing a front moving into fluid in equilibrium along the direction $\hat{\mathbf{k}}$. The solution depends on time and space only through the variable $\xi=\hat{\mathbf{k}}\cdot\mathbf{x}-vt$, where $v$ is the front velocity. Let the front position be $\xi=\xi_{0}$. At this point, the $C^{\alpha}$’s are continuous but the lateral $\xi-$ derivatives ${C^{\prime}}^{\alpha}_{+}$ and ${C^{\prime}}^{\alpha}_{-}$ are different. Therefore, taking the difference of the eqs. (89) in front of and behind the front, the terms from the collision integral cancel and we get
with a prime denoting a $\xi$- derivative.
It is clear that eqs. (90) are the same as eqs. (91) under the identification $\omega=vk$, $\mathbf{k}=k\hat{\mathbf{k}}$ and in the free streaming limit where the collision term is neglected. We shall take advantage of this simplification and not consider terms coming from the collision term in the following. This is not an approximation but rather the definition of the propagation speeds.
VI.2 The complete set of equations of motion
In the linearized approximation
The free streaming equations of motion are then energy-momentum conservation
and the moment equations
$\left\langle X^{\beta}p^{\mu}\left(\beta_{\nu,\mu}p^{\nu}+\sum_{\alpha=1}^{4}C_{\alpha,\mu}X^{\alpha}\right)\right\rangle=0$ (94)
$\beta=1-4$, or else, working in the rest frame where $u^{\mu}=(1,0,0,0)$ and setting $t=1$
$\left\langle p^{0}p^{\mu}p^{\nu}\right\rangle\beta_{\nu,\mu}+\sum_{\alpha=1}^{4}\left[\dot{C}_{\alpha}\left\langle p^{02}X^{\alpha}\right\rangle+C_{\alpha,i}\left\langle p^{i}p^{0}X^{\alpha}\right\rangle\right]$ $=$ 00 0
$\left\langle X^{\beta}p^{\mu}p^{\nu}\right\rangle\beta_{\nu,\mu}+\sum_{\alpha=1}^{4}\left[\dot{C}_{\alpha}\left\langle X^{\beta}p^{0}X^{\alpha}\right\rangle+C_{\alpha,i}\left\langle X^{\beta}p^{i}X^{\alpha}\right\rangle\right]$ $=$ 00 0 (95)
The constraint eq. (22) implies that $\left\langle p^{02}X^{\alpha}\right\rangle=\left\langle p^{0}p^{j}X^{\alpha}\right\rangle=0$, so these equations simplify to
$\left\langle p^{0}p^{\mu}p^{\nu}\right\rangle\beta_{\nu,\mu}$ $=$ 00 0 (96)
Computing the averages we find the full set of equations in the free-streaming regime
$\frac{\dot{t}}{t}+\frac{1}{3}u^{j}_{,j}$ $=$ 00 0
The different coefficients in the equations are
$A$ $=$ $\left[\Theta_{5}^{2}-\gamma\Theta_{4}^{1}\right]$ (100)
$F$ $=$ $\left[D_{2}\Theta_{4}^{2}-\frac{\Theta^{1}_{4}}{\Theta^{0}_{3}}\Theta_{5}^{2}+\Theta_{6}^{3}-D_{3}\Theta_{4}^{1}\right]$ (101)
$G$ $=$ $\left[\Theta_{7}^{4}-\gamma\Theta_{6}^{3}\right]$ (102)
$K$ $=$ $\left[D_{2}\Theta_{5}^{3}-\frac{\Theta^{1}_{4}}{\Theta^{0}_{3}}\Theta_{6}^{3}+\Theta_{7}^{4}-D_{3}\Theta_{5}^{2}-\gamma F\right]$ (103)
$M$ $=$ $\left[D_{2}^{2}\Theta_{3}^{2}+\left(\frac{\Theta^{1}_{4}}{\Theta^{0}_{3}}\right)^{2}\Theta_{5}^{2}+\Theta_{7}^{4}+D_{3}^{2}\Theta_{3}^{0}-2D_{2}\frac{\Theta_{4}^{1}}{\Theta_{3}^{0}}\Theta_{4}^{2}-2\frac{\Theta_{4}^{1}}{\Theta_{3}^{0}}\Theta_{6}^{3}+2D_{2}\Theta_{5}^{3}\right.$ (104)
-- $\left.2D_{2}D_{3}\Theta_{3}^{1}+2D_{3}\frac{\Theta_{4}^{1}}{\Theta_{3}^{0}}\Theta_{4}^{1}-2D_{3}\Theta_{5}^{2}\right]$
where $D_{2}$, $D_{3}$ and $\gamma$ are defined in eqs. (78), (83) and (85), respectively. Note that all these coefficients vanish in the Anderson-Witting RTA.
VI.3 SVT Decomposition
As it is well known, the equations of motion can be further decoupled by decomposing the deviations from equilibrium into scalar, vector and tensor quantities. We write
$u^{i}$ $=$ $\vartheta^{i}+\nabla_{i}\vartheta$
$C_{4i}$ $=$ $c_{4i}+\nabla_{i}c_{4}$
$C_{1ij}$ $=$ $\left(\nabla_{i}\nabla_{j}-\frac{1}{3}\Delta_{ij}\nabla^{2}\right)c_{1}+c_{1i,j}+c_{1j,i}+c_{1ij}$
$C_{3ij}$ $=$ $\left(\nabla_{i}\nabla_{j}-\frac{1}{3}\Delta_{ij}\nabla^{2}\right)c_{3}+c_{3i,j}+c_{3j,i}+c_{3ij}$
$+$ $\nabla_{i}c_{2jk}+\nabla_{j}c_{3ik}+\nabla_{k}c_{2ij}+c_{2ijk}$
We shall discuss in some detail the simplest tensor sector, and give the results for the vector and scalar ones, which follow the same structure. In the free streaming limit the equation for $c_{2ijk}$ reduces to $\dot{c}_{2ijk}=0$ and we shall not discuss it.
VI.3.1 Tensor Sector
Considering only tensor quantities in eqs. (99) we get
$\Theta_{5}^{2}\dot{C}_{1ij}+\frac{3}{7}\Theta_{6}^{3}\nabla^{2}C_{2ij}$ $=$ 00 0 (106)
$\Theta_{6}^{3}C_{1ij}+3\Theta_{7}^{4}\dot{C}_{2ij}+GC_{3ij}$ $=$ 00 0 (107)
$G[\frac{3}{7}\nabla^{2}C_{2ij}+\dot{C}_{3ij}]$ $=$ 00 0 (108)
Explicitly
$03G\Theta^{2}_{5}\Theta^{4}_{7}(i\omega)[\omega^{2}-\frac{1}{7}k^{2}]=$ 0 (110)
Observe that although the determinant in eq. (109) becomes singular for the Anderson-Witting RTA, where $G=0$, the dispersion relation is well defined there, and yields the propagation speeds $1/\sqrt{7}$ and 00, independently of the choice of the function $F$.
As we shall show presently, the propagation speeds depend on the choice of the function $F$ in the vector and scalar sectors.
VI.3.2 Vector Sector
The vectorial terms in eq. (99) are
$\Theta_{3}^{0}\dot{\vartheta}^{i}+\frac{2}{5}\Theta_{4}^{1}\nabla^{2}c_{1i}+\frac{2}{5}A\nabla^{2}c_{3i}$ $=$ 00 0 (111)
$\nabla_{j}\left[\Theta_{4}^{1}\vartheta_{i}+2\Theta_{5}^{2}\dot{c}_{1i}+\frac{24}{35}\Theta_{6}^{3}\nabla^{2}c_{2i}+Fc_{4i}\right]$ $=$ 00 0 (112)
$\left(\nabla_{j}\nabla_{k}-\frac{1}{5}\Delta_{jk}\nabla^{2}\right)\left[\frac{2}{3}\Theta_{6}^{3}c_{1i}+\Theta_{7}^{4}\dot{c}_{2i}+\frac{2}{3}Gc_{3i}\right]$ $=$ 00 0 (113)
$\nabla_{j}\left[A\vartheta_{i}+\frac{24}{35}G\nabla^{2}c_{2i}+2G\dot{c}_{3i}+Kc_{4i}\right]$ $=$ 00 0 (114)
$\frac{2}{5}F\nabla^{2}c_{1i}+\frac{2}{5}K\nabla^{2}c_{3i}+M\dot{c}_{4i}$ $=$ 00 0 (115)
Wherefrom we get the caractheristic equation
VI.3.3 Scalar Sector
Keeping only scalar terms in eq. (99) we get
$\dot{t}+\frac{1}{3}\nabla^{2}\vartheta$ $=$ 00 0 (117)
$\nabla_{i}\left[\Theta_{3}^{0}\left(t+\dot{\vartheta}\right)+\frac{4}{15}\Theta_{4}^{1}\nabla^{2}c_{1}+\frac{4}{15}A\nabla^{2}c_{3}\right]$ $=$ 00 0 (118)
$2\left(\nabla_{i}\nabla_{j}-\frac{1}{3}\Delta_{ij}\nabla^{2}\right)\left[\Theta_{4}^{1}\vartheta+\Theta_{5}^{2}\dot{c}_{1}+\frac{9}{35}\Theta_{6}^{3}\nabla^{2}c_{2}+Fc_{4}\right]$ $=$ 00 0 (119)
$6\left[\nabla_{i}\nabla_{j}\nabla_{k}-\frac{1}{5}\nabla^{2}\Delta_{ij}\nabla_{k}-\frac{1}{5}\nabla^{2}\Delta_{ik}\nabla_{j}-\frac{1}{5}\nabla^{2}\Delta_{jk}\nabla_{i}\right]\left[\Theta_{6}^{3}c_{1}+\Theta_{7}^{4}\dot{c}_{2}+Gc_{3}\right]$ $=$ 00 0 (120)
$2\left(\nabla_{i}\nabla_{j}-\frac{1}{3}\Delta_{ij}\nabla^{2}\right)\left[A\vartheta+\frac{9}{35}G\nabla^{2}c_{2}+G\dot{c}_{3}+Kc_{4}\right]$ $=$ 00 0 (121)
Leading to the dispersion relations
VI.4 Results
$F[x]=x^{a}$ (124)
For which
$\Theta^{n}_{m}=\Gamma[2+m-an]$ (125)
To avoid infrared divergences in the equations of motion we require $a\leq 2$. This family includes the Marle and Anderson-Witting RTAs as particular cases, namely $a=0$ and $a=1$, respectively.
Finally, in Fig. (3) we plotted the curves that correspond to the speed of the fastest mode of each sector. The top dotted horizontal line corresponds to the AW ($a=1$) speed, short-dashed line immediately below corresponds to the velocities of the scalar fastest mode of our model. Middle dotted horizontal line and dot-dashed middle curve correspond respectively to the vector mode speed for AW ($a=1$) and to the speeds of our model. Bottom long-dashed horizontal line is the speed of the tensor mode, which agrees with the AW speed over the entire interval of $a$ values considered. They verify $v_{T}<v_{V}<v_{S}$, as expected [39].
Figure 1: (Color online) Speeds of the two vector modes from eq. (116). The fastest mode (top dot-dashed blue curve) attains the same maximum speed for $a=1$ and for $a\to-\infty$ (top dotted light-blue horizontal line), indicating that the AW value is not exceeded at any value of $a$. Observe that the slowest mode speed (bottom, dashed light blue curve) is zero for $a=1$, so we recover the AW case [47] where only one non-null mode exists.
Figure 2: (Color online) Speeds of the three scalar modes from eq. (123). We see that the maximum speed of the fastest mode (top red short-dashed curve) corresponds to the AW solution $a=1$ (top horizontal orange dotted line). The intermediate speed mode (long-dashed orange curve in the middle of the figure) attains its minimum value also at the AW value $a=1$ (bottom horizontal yellow dotted line), and the speeds of this mode never exceed the ones of the fastest mode. These two modes are the generalization of the AW modes found elsewhere. The bottom, single-line purple curve corresponds to the speeds of a new, slowest mode, whose velocity for $a=1$ is zero. Thus we see that the AW case, for which there are only two non-null modes, is consistently included in our formalism.
Figure 3: (Color online) Comparison of the maximum propagation speeds of each sector. Top dotted horizontal line corresponds to the AW ($a=1$) scalar mode speed, short-dashed line immediately below corresponds to the velocities of the scalar fastest mode of our model. Middle dotted horizontal line and dot-dashed middle curve correspond respectively to the vector mode speed for AW ($a=1$) and to the speeds of our model. Bottom long-dashed horizontal line is the speed of the tensor mode, which agrees with the AW speed over the entire interval of $a$ values considered. They verify $v_{T}<v_{V}<v_{S}$.
VII Conclusions
The aim of this work was to compute the propagation speeds for a conformal relativistic fluid, assuming hydrodynamics is derived by taking moments of a RTA kinetic equation where the relaxation time becomes a function of momentum, $\tau=\tau_{0}/F[-\beta_{\mu}p^{\mu}]$, suitably modified to make it consistent with energy-momentum conservation. We showed that it is possible to build a kinetic equation within this family which matches any given spectrum for the full Boltzmann equation.
The derivation of hydrodynamics involves assuming a particular parameterization for the 1pdf. We chose the criteria that the parameterized 1pdf should contain the second order Chapman-Enskog solution of the generalized RTA kinetic equation as a special case. Using the moments method we then obtained the sets of the general linear hydrodynamic equations for scalar, vector and tensor perturbations. The choice of which moments to take is determined by the requirement that the dynamics of the perameters enforces positive entropy production.
To analyze a concrete case, we specialized the general equations to the case where $F=\left(\beta_{\mu}p^{\mu}\right)^{a}$, which includes the choices of Anderson-Witting $a=1$ [20, 21] and Marle $a=0$ [18, 19], widely used in the literature, as particular cases. The interpolating values $0\leq a\leq 1$ have already been discussed in the literature [22, 23], motivated by its possible application to improve the description of relativistic heavy ion collisions. Here we have included the full range $-\infty\leq a\leq 2$, which is maximal because larger values of $a$ leads to infrared divergences in the coefficients of the hydrodynamic equations. Allowing for a negative $a$ allows us to explore distributions heavily biased towards hard modes.
We obtained the linearized propagation equations for the tensor, vector and scalar modes and compute the corresponding velocities. The propagation speed is the phase velocity for plane waves in the free streaming limit. Note that though we included the $F$ function through the collision term, taking the free streaming limit does not make it irrelevant, because the chosen parameterization of the 1pdf, dictated by matching to the Chapman-Enskog solution, depends explicitly on $F$.
At the considered order of the 1-PDF the number of tensor, vector and scalar modes are respectively 2, 5 and 6.
For the tensor modes the propagation speed is actually independent of $a$ and agrees with the free streaming Anderson-Witting value [47].
In the vector sector, besides the trivial solution $v_{v}=0$, there are two propagation speeds shown in Fig. (1). There we saw that the fastest propagation speed is bounded above by the Anderson-Witting value [47], which is reached at $a=1$. The slower mode has $v_{v}=0$ for $a=1$. Therefore the number of dynamical vector modes in the Anderson-Witting limit reduces to $3$, as it must.
For the scalar sector we obtained three different propagation speeds as is shown in Fig. (2). As in the tensor and vector sectors, the fastest mode has maximum velocity at $a=1$, where we recover the free streaming Anderson-Witting result [47]. For the intermediate mode we also recover the lower AW speed when $a=1$. The speeds of the slowest mode are significantly lower than the other two scalar modes, and vanishes for $a=1$. In this way we recover the right number od dynamical scalar modes ($4$) in the Anderson-Witting case.
In Fig. (3) we compared the speeds of the three fastest modes. We see that, as expected, they satisfy $v_{T}<v_{v}<v_{s}$ throughout the whole range of $a$ values [39].
We believe that the main contributions of this work are that the theory remains causal for the full range of values of $a$, and that actually the fastest propagation speeds are found in the Anderson-Witting limit $a=1$. Causality is expected because, as we have already said, the theory is built to enforce thermodynamic stability, and it is known that stability, causality and covariance are closely linked [48, 49, 50, 51, 52, 53].
To the best of our knowledge, the fact that the Anderson-Witting RTA [20, 21] produces the fastest propagation speeds is a new result. This has deep implications for the description of strong shocks in relativistic fluids [55], which we expect to elaborate on in a separate contribution.
Acknowledgements.

### v2

@@ A.2 Second order Ch-En solution
Enforcing the constraint eq. (27) yields eq. (60).
Referencias
  • [1] P. Romatschke and U. Romatschke, Relativistic Fluid Dynamics In and Out of Equilibrium And Applications to Relativistic Nuclear Collisions, Cambridge Univ. Press, Cambridge, UK (2019).
  • [2] E. Calzetta, Real Relativistic Fluids in Heavy Ion Collisions, in Geometric, Algebraic and Topological Methods for Quantum Field Theory; L. Cano, A. Cardona, H. Ocampo, A. F. R. Lega Eds; World Scientific, Singapore, p 155 (2016).
  • [3] N. Mirón-Granese, E. Calzetta and A Kandus, Primordial Weibel instability, JCAP 2022, 028 (2022).
  • [4] N. Mirón-Granese and E. Calzetta, Primordial Gravitational Waves Amplification from Causal Fluids, Phys. Rev. D 97, 023517 (2018).
  • [5] N. Mirón-Granese, Relativistic Viscous Effects on the Primordial Gravitational Waves Spectrum JCAP 2021, 008 (2021).
  • [6] E. Calzetta and B. L. Hu, Nonequilibrium Quantum Field Theory, Cambridge Univ. Press, Cambridge, UK(2008).
  • [7] L. Rezzolla and O. Zanotti, Relativistic Hydrodynamics (Oxford University Press, Oxford, 2013).
  • [8] G. Denicol and D. Rischke, Microscopic Foundations of Relativistic Fluid Dynamics (Springer, Berlin, 2021). f
  • [9] W. Israel, Nonstationary irreversible thermodynamics: A causal relativistic theory, Ann. Phys. (NY) 100, 310 (1976).
  • [10] I. S. Liu, I. Müller and T. Ruggeri, Relativistic thermodynamics of gases, Ann. Phys. 169, 191 (1986).
  • [11] R. Geroch and L. Lindblom, Dissipative relativistic fluid theories of divergence type, Phys. Rev. D 41, 1855 (1990).
  • [12] J. Peralta-Ramos and E. Calzetta, Divergence-type nonlinear conformal hydrodynamics, Phys. Rev. D 80, 126002 (2009).
  • [13] P. L. Bhatnagar, E. P. Gross, M. Krook. A Model for Collision Processes in Gases. I. Small Amplitude Processes in Charged and Neutral One-Component Systems, Phys. Rev. 94, 511 (1954)
  • [14] C. Marle, Sur l’etabissement des équations de l’hydrodynamique des fluids relativistes dissipatifs. I. - L’équation de Boltzmann relativiste., Ann. Inst. Henri Poincaré (A), 10, 67 (1969).
  • [15] C. Marle, Sur l’etabissement des équations de l’hydrodynamique des fluids relativistes dissipatifs. II. - Méthodes de résolution approchée de l’equation de Boltzmann relativiste., Ann. Inst. Henri Poincaré (A), 10, 127 (1969).
  • [16] J. L. Anderson and H. R. Witting, A relativistic Relaxation-Time Model for the Boltzmann Equation, Physica 74, 466 (1974)
  • [17] J. L. Anderson and H. R. Witting, Relativistic Quantum Transport Coefficients, Physica 74, 489 (1974).
  • [18] G. S. Denicol and J. Noronha, Spectrum of the Boltzmann collision operator for $\lambda\phi^{4}$ theory in the classical regime, Phys. Lett. B 850, 138487 (2024).
  • [19] G. S. Rocha, C. V. P. de Brito and G. S. Denicol, Hydrodynamic theories for a system of weakly self-interacting classical ultra-relativistic scalar particles: microscopic derivations and attractors, Phys. Rev. D 108, 036017 (2023).
  • [20] M. Dudynski, Spectral properties of the linearized Boltzmann operator in Lp for $1\leq p\leq\infty$, J. Stat. Phys. 153, 1084 (2013).
  • [21] L. Luo and H. Yu, Spectrum analysis of the linearized relativistic Landau equation, J. Stat. Phys. 163, 914 (2016).
  • [22] J. Hu, Relaxation time approximation revisited and pole/cut structure in retarded correlators, e-print arXiv:2409.05131 (2024).
  • [23] K. Dusling, G. D. Moore, and D. Teaney, Radiative energy loss and v 2 spectra for viscous hydrodynamics, Phys. Rev. C 81, 034907 (2010).
  • [24] M. Luzum and J.-Y. Ollitrault, Constraining the viscous freeze-out distribution function with data obtained at the BNL Relativistic Heavy Ion Collider (RHIC), Phys. Rev. C 82, 014906 (2010)
  • [25] G. Rocha, G. Denicol and J. Noronha, Novel Relaxation Time Approximation to the Relativistic Boltzmann Equation , Phys. Rev. Lett. 127, 042301 (2021).
  • [26] G. Rocha, M. Ferreira, G. Denicol and J. Noronha, Transport coefficients of quasiparticle models within a new relaxation time approximation of the Boltzmann equation, Phys. Rev. D 106, 036022 (2022).
  • [27] E. Calzetta and J. Peralta-Ramos, Linking the hydrodynamic and kinetic description of a dissipative relativistic conformal theory, Phys. Rev. D 82, 106003 (2010).
  • [28] J. Peralta-Ramos and E. Calzetta, Macroscopic approximation to relativistic kinetic theory from a nonlinear closure, Phys. Rev. D 87, 034003 (2013).
  • [29] A. Kurkela and U. Wiedemann, Analytic structure of nonhydrodynamic modes in kinetic theory , Eur. Phys. J. C 79, 776 (2019).
  • [30] G. Wilka and Z. Włodarczyk, Beyond the relaxation time approximation, Eur. Phys. J. A57, 221 (2021).
  • [31] S. Mitra, Relativistic hydrodynamics with momentum-dependent relaxation time , Phys. Rev. C 103, 014905 (2021).
  • [32] S. Mitra, Correspondence between momentum-dependent relaxation time and field redefinition of relativistic hydrodynamic theory, Phys. Rev. C 105, 014902 (2022).
  • [33] V. Ambruş and E. Molnár, Shakhov-type extension of the relaxation time approximation in relativistic kinetic theory and second-order fluid dynamics, e-print arXiv:2311.11603.
  • [34] V. Ambruş and E. Molnár, High-order Shakhov-like extension of the relaxation time approximation in relativistic kinetic theory, e-print arXiv:2401.04017.
  • [35] J. Hu, Full-order mode analysis within a mutilated relaxation time approximation, eprint arXiv:2310.05606
  • [36] D. Dash, S. Bhadury, S. Jaiswal and A. Jaiswal, Extended relaxation time approximation and relativistic dissipative hydrodynamics, Phys. Lett. B, 831, 13720 (2022).
  • [37] D. Dash, S. Bhadury, S. Jaiswal and A. Jaiswal, Relativistic second-order viscous hydrodynamics from kinetic theory with extended relaxation-time approximation, Phys. Rev. C 108, 064913 (2023).
  • [38] T. Bhattacharyya, Non-extensive Boltzmann Transport Equation: the Relaxation Time Approximation and Beyond, Physica A 624, 128910 (2023).
  • [39] G. Rocha, D. Wagner, G. Denicol, J. Noronha, and D. Rischke, D.H. Theories of Relativistic Dissipative Fluid Dynamics. Entropy 26, 189 (2024).
  • [40] M. Strickland, Hydrodynamization and resummed viscous hydrodynamics, arXiv:2402.09571, Contributed chapter to Quark-Gluon Plasma 6 book
  • [41] L. Gavassino, Mapping GENERIC Hydrodynamics into Carter’s Multifluid Theory. Symmetry 16, 78 (2024).
  • [42] D. Wagner and L. Gavassino, The regime of applicability of Israel-Stewart hydrodynamics, Phys. Rev. D 109, 016019 (2024).
  • [43] L. Gavassino, M. M. Disconzi and J. Noronha, Universality Classes of Relativistic Fluid Dynamics: Foundations, Phys. Rev. Lett. 132, 222302 (2024).
  • [44] L. Gavassino, M. M. Disconzi and J. Noronha, Universality classes of relativistic fluid dynamics: Applications, Phys. Rev. D 109, 096041 (2024).
  • [45] S. Chapman and T. G. Cowling, The Mathematical Theory of Non-Uniform Gases, 3rd. ed. ,Cambridge Univ. Press, Cambridge (1970).
  • [46] G. S. Denicol, H. Niemi, E. Molnár and D. H. Rischke, Derivation of transient relativistic fluid dynamics from the Boltzmann equation, Phys. Rev. D 85, 11047 (2012).
  • [47] D. Wagner, A. Palermo and V. Ambrus, Inverse-Reynolds-dominance approach to transient fluid dynamics, Phys. Rev. D 106, 016013 (2022).
  • [48] S. Bhadury, M. Kurian, V. Chandra and A. Jaiswal, First order dissipative hydrodynamics and viscous corrections to the entropy four-current from an effective covariant kinetic theory, J. Phys. G 47 (2020) 8, 085108
  • [49] S. Bhadury, M. Kurian, V. Chandra and A. Jaiswal, Second order relativistic viscous hydrodynamics within an effective description of hot QCD medium, J. Phys. G: 48, 105104, (2021).
  • [50] S. Diles, A. Miranda, L. Mamani, A. Echemendia and V. Zanchin, Third-order relativistic fluid dynamics at finite density in a general hydrodynamic frame, arXiv:2311.01232.
  • [51] G. Perna and E. Calzetta, Linearized dispersion relations in viscous relativistic hydrodynamics, Phys. Rev. D 104, 096005 (2021).
  • [52] L. Cantarutti and E. Calzetta, Dissipative-type theories for Bjorken and Gubser flows, Int. J. Mod. Phys. A 35, 2050074 (2020).
  • [53] M. Aguilar and E. Calzetta, Causal relativistic hydrodynamics of conformal Fermi-Dirac gases, Phys, Rev. D 95, 076022 (2017)
  • [54] H. Struchtrup, Macroscopic Transport Equations for Rarefied Gas Flows (Springer, Berlin, 2005).
  • [55] G. Boillat and T. Ruggeri, On the evolution law of weak discontinuities for hyperbolic quasi-linear systems, Wave Motion 1, 149 (1979).
  • [56] G. Boillat and T. Ruggeri, Hyperbolic principal subsystems: entropy convexity and subcharacteristic conditions, Arch. Rational Mech. Anal. 137, 305 (1997).
  • [57] G. Boillat and T. Ruggeri, Moment equations in the kinetic theory of gases and wave velocities, Continuum Mech. Thermodyn. 9, 205 (1997).
  • [58] G. Boillat and T. Ruggeri, Relativistic gas: Moment equations and maximum wave velocity, J. Math. Phys. (N.Y.) 40, 6399 (1999).
  • [59] I. Müller, Speeds of propagation in classical and relativistic extended thermodynamics, Living Rev. Relativity 2, 1 (1999).
  • [60] S. Mitra, Causality and stability analysis of first-order field redefinition in relativistic hydrodynamics from kinetic theory, Phys. Rev. C 105, 054910 (2022).
  • [61] M. Heller, A. Serantes, M. Spaliński and B. Withers, The Hydrohedron: Bootstrapping Relativistic Hydrodynamics, arXiv:2305.07703v2
  • [62] N. Abboud, E. Speranza and J. Noronha, Causal and stable first-order chiral hydrodynamics , arXiv:2308.02928.
  • [63] C. de Brito, G. Rocha and G. Denicol, Hydrodynamic theories for a system of weakly self-interacting classical ultra-relativistic scalar particles: causality and stability , arXiv:2311.07272.
  • [64] R. Hoult and P. Kovtun, Causality and classical dispersion relations, Phys. Rev. D 109 (2024) 4, 046018.
  • [65] D-L Wang and S. Pu, Stability and causality criteria in linear mode analysis: stability means causality, Phys. Rev. D 109, L031504 (2024).
  • [66] L. Gavassino, M. Disconzi and J. Noronha, Dispersion relations alone cannot guarantee causality, Phys. Rev. Lett. 132, 162301 (2024).
  • [67] W. Israel, Relativistic theory of shock waves, Proc. R. Soc. A 259, 129 (1960).
  • [68] E. Calzetta, Steady asymptotic equilibria in conformal relativistic fluids, Phys. Rev. D 105, 036013 (2022).
  • [69] W. Israel, Covariant fluid mechanics and thermodynamics: An introduction, in A. M. Anile and Y. Choquet-Bruhat (eds.) Relativistic Fluid Dynamics(Springer, New York, 1988), p. 152.
  • [70] G. Moore, Hydrodynamics as $v_{s}\to c$, arXiv:2404.01968v2 [hep-ph] (2024).
  • [71] W. Israel, The relativistic Boltzmann equation. In: L. O’Raifeartaigh (ed.), General Relativity: Papers in Honour of J. L. Synge, pp. 201–241. Clarendon, Oxford (1972).
  • [72] S. R. de Groot, W. A. van Leeuwen, and C. G. van Weert, Relativistic Kinetic Theory (North-Holland, Amsterdam, 1980).
  • [73] J. M. Stewart, Non-Equilibrium Relativistic Kinetic Theory (Springer, New York, 1971).
  • [74] R. Liboff, Kinetic Theory (Springer, New York, 2003).
  • [75] C. Cercignani and G. Medeiros Kremer, The Relativistic Boltzmann Equation: Theory and Applications (Birkhauser, Basel, 2002).
  • [76] C. Eckart, The Thermodynamics of Irreversible Processes. III. Relativistic Theory of the Simple Fluid, Phys. Rev. 58, 919 (1940)
  • [77] L. D. Landau and E. M. Lifshitz, Fluid Mechanics, Pergamon Press, Oxford (1959).
  • [78] T. Dore,, L. Gavassino, D. Montenegro, M. Shokri and G. Torrieri, Fluctuating relativistic dissipative hydrodynamics as a gauge theory, Ann. of Phys. 442, 168902 (2022).
  • [79] S. Bhattacharyya, S. Mitra, S. Roy and R. Singh, Field redefinition and its impact in relativistic hydrodynamics, arXiv:2409.15387 [nucl-th] (2024).
  • [80] S. Bhattacharyya, S. Mitra and S. Roy, Causality and Stability in relativistic hydrodynamic theory - a choice to be endured, Phys. Lett. B 856, 138918 (2024).
  • [81] W. Hiscock and L. Lindblom, Stability and Causality in Dissipative Relativistic Fluids, Ann. Phys. (NY) 151, 466 (1983).
  • [82] W. Israel and M. Stewart, Transient relativistic thermodynamics and kinetic theory, Ann. Phys. (NY) 118, 341 (1979).